# Fantastic Weights and How to Find Them: Where to Prune in Dynamic Sparse Training

### v1

@@ Appendix H Global versus Local Pruning
In our study we have decided to focus on local pruning as this is the typical approach in the DST literature (e.g., RigL [12] (Algorithm 1), SNFS [11] (Figure 1, where pruning and redistribution occur layer-wise), SET [40]). In contrast, global pruning computes and compares the importance scores associated with the pruning criterion with all the parameters in the network. Hence, global pruning may change the sparsity distributions in each layer in an unsupervised manner. It is important to note that global pruning may also have one small disadvantage: It may carry a higher risk of disconnecting the model if the importance scores $s(\theta^{l}_{i,j})$ are not balanced across different layers. Using local pruning may avoid this risk and omits the need for investigating layer-wise normalization schemes.
@@
However, we do consider the question of how the studied pruning criteria perform when paired with global pruning an interesting one. We perform a small experiment on ResNet-56 with CIFAR10 in which we compare the different pruning criteria in the global- and local-pruning setup on densities $0.05$, $0.1$, and $0.15$. We present the results in Figure 13.171717We leave the $\mathcal{C}_{\mathrm{SET}}$ out due to its high similarity to the magnitude pruning. We observe that the results obtained in both global- and local- setups are indeed similar, with a slight preference for global pruning. By analyzing the plots, we can see that in this case, the ranking results of the pruning criteria from the main text are indeed preserved. We consider the subject of using global pruning a very intriguing topic to investigate further in future research.
@@
Appendix I Additional Jaccard Index Plots
@@
Figure 16: The performance of the pruning criteria with respect to the density on different datasets. Each row corresponds to a different model-dataset pair. The first column corresponds to the random growth, the second one to the gradient growth. Please note the logarithmic scale in the x-axis. We may observe that the differences between the methods become more visible predominantly in the low-density regime. DST almost always outperforms the static sparse training (cyan line).
Appendix J Learning Curves
@@
Figure 19: The validation accuracy versus the training epoch for different pruning criteria and densities on the small-CNN (CIFAR10 dataset).

### v2

@@ Appendix I Global versus Local Pruning
In our study, we have decided to focus on local pruning as this is the typical approach in the DST literature. Examples include RigL [12] (Algorithm 1), SNFS [11] (Figure 1, where pruning and redistribution occur layer-wise), SET [42]. In contrast, global pruning computes and compares the importance scores associated with the pruning criterion with all the parameters in the network. Hence, global pruning may change the sparsity distributions in each layer in an unsupervised manner. Note that global pruning may also have one small disadvantage: It may carry a higher risk of disconnecting the model if the importance scores $s(\theta^{l}_{i,j})$ are not balanced across different layers. Using local pruning may avoid this risk and omit the need for investigating layer-wise normalization schemes.
@@
However, we do consider the question of how the studied pruning criteria perform when paired with global pruning an interesting one. We perform a small experiment on ResNet-56 with CIFAR10 in which we compare the different pruning criteria in the global- and local-pruning setup on densities $0.05$, $0.1$, and $0.15$. We present the results in Figure 15.181818We leave the $\mathcal{C}_{\mathrm{SET}}$ out due to its high similarity to the magnitude pruning. We observe that the results obtained in both global- and local- setups are indeed similar, with a slight preference for global pruning. By analyzing the plots, we can see that, in this case, the ranking results of the pruning criteria from the main text are indeed preserved. We consider the subject of using global pruning a very intriguing topic to investigate further in future research.
@@
Appendix J Effect of Batch Size and Pruning Schedule on Update Period
When considered together with the batch size and dataset size, the update period can be interpreted as the number of data samples the model sees (and trains on) before performing the mask update. In consequence, one may expect that with the increase of the batch size, the best choice of the update period will shift to lower values, keeping the total number of seen examples similar. To visualize this effect, we run a comparison on the MLP-CIFAR10 setup and observe just this behavior (see Figure 16(c) and Figure 16(d)).
Another potential confounding factor is the pruning rate schedule. In all our main experiments, we used the cosine schedule, which is known to perform well in the DST setup [12]. In this section, we additionally test this schedule on the MLP-CIFAR10 setup against two different schedulers: linear and constant. The linear schedule multiplies every 600 optimization steps the current decay rate by a constant factor (0.99). The constant schedule keeps the prune fraction fixed at 0.5 throughout the whole training. The results are presented in Figure 16(a) and Figure 16(b). We observe that the choice of the pruning factor schedule does not significantly influence the best value of the update period.
Figure 16: Additional results for the sensitivity to hyperparameters in the update-period study. We perform the study on random growth. From the left, we plot the results for (a) linear scheduler and default (i.e., 128) batch size (BS), (b) constant scheduler and default batch size, (c) cosine scheduler and default batch size (the same setup as in main paper),(d) cosine scheduler and batch size 256.
Appendix K Additional Jaccard Index Plots
@@
Moreover, we analyze the statistics of the never-removed connections. To be precise, we consider the masks at the end of training for the MLP, CNN, and ResNet-56 models on CIFAR10. Next, we measure the number of weights that are always retained by every pruning criterion for the same seed and we divide it by the number of all remaining (unpruned) weights in the model. We do the same on the negation of the masks to get the number of the always removed weights and divide it by the number of all removed weights of the model. We present the results in Figure 19. For the MLP model, around $23\%$ of all the retained weights are shared across all the models. For CNN, this number is higher and goes up to $37\%$, while for the ResNet-56 it is equal to $11\%$. In general, the smaller overlap of the pruning methods for the ResNet model is also something that we observe in the main paper in Figure 6b.
Finally, we report the standard deviations for the Jaccard Index plots from the main text in Figures 20 and 21. Note that the mean Jaccard Index for the sets chosen for removal in the main text was computed by first taking a mean over different seeds for each layer and then averaging the results over all layers. Hence, in Figure 20, the standard deviation is calculated for each layer separately and then averaged. For the end-similarity, the Jaccard Index was computed by first taking the mean over all layers for a given seed and then taking the mean over all random seeds. Hence, in Figure 21, we report the standard deviation for averaging over the different random seeds.
@@
Figure 19: The number of weights retained by all the pruning criteria relative to the number of all unpruned weights (in blue). The number of weights removed by all the pruning criteria relative to the number of all pruned weights (in pink).
Figure 20: The mean standard deviation computed for the Jaccard index between the sets of weights chosen for removal during the first update of the sparse connectivity.
Figure 21: The standard deviation of the same index computed between the masks obtained by different pruning criteria at the end of training. The rows and columns represent the pruning criteria between which the index is computed.
Figure 22: The performance of the pruning criteria with respect to the density on different datasets. Each row corresponds to a different model-dataset pair. The first column corresponds to the random growth, the second one corresponds to the gradient growth. Please note the logarithmic scale in the x-axis. We may observe that the differences between the methods become more visible predominantly in the low-density regime. DST almost always outperforms the static sparse training (cyan line).
Appendix L Learning Curves
@@
Figure 25: The validation accuracy versus the training epoch for different pruning criteria and densities on the small CNN (CIFAR10 dataset).
@@
wDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">